# The SRG/eROSITA all-sky survey

### v1

@@ 3.1 X-ray surface brightness distribution
The sector 270–360 deg in Fig. 10 shows a bump starting around , likely related to the Fornax A group. In Fig. 11 we then show the surface brightness in a small sector towards and beyond the Fornax A group. It is clear that significant group emission spans at least 1 deg. Moreover, the indication for an apparent emission Bridge between the Fornax cluster and the Fornax A group is confirmed also in the surface brightness profile.
@@
To get a better handle on the origin of these surface brightness extensions and excess regions, we compare to other tracers of the large-scale structure in and around the Fornax cluster in the next sections.
@@ 3.2 Distribution of cluster member galaxies
The most obvious additional tracer of structure is the distribution of galaxies in and around the Fornax cluster. As the Fornax cluster redshift is 0.005, we select all objects classified as galaxies from the NASA/IPAC Extragalactic Database (NED)444https://ned.ipac.caltech.edu in the image area with a redshift less than 0.01. All of these are assumed to be member galaxies. With this selection we aim for high completeness of member galaxies. Since velocity dispersion measurements for the Fornax cluster reach km/s and above (e.g., Drinkwater et al., 2001; Pota et al., 2018) we want to ensure to include all galaxies within at least km/s, which corresponds to a range of at least –. Figure 12 shows the member galaxy distribution.
In the central region up to about , the galaxy distribution exhibits an elongation in the east-west direction. On larger scales, well beyond , one can notice a correlation of X-ray surface brightness excess regions and galaxy density. In particular, the Finger 1 and 2, and Excess 1, 2, and 3 seem to have their counterpart in the galaxy distribution. Moreover, also the X-ray emission Bridge towards the NGC 1316 group has an excess of galaxies; related to the lopsided distribution of Fornax A group dwarf galaxies towards the north-east discussed in Sect. 1 and (Su et al., 2021). The underlying reason for this offset between galaxy distribution and Fornax A intragroup gas might be related to the collisional gas lagging behind the collisionless galaxies in their common motion, e.g., due to ram pressure relative to the intracluster medium of the Fornax cluster outskirts; however, a more detailed investigation of this scenario is beyond the scope of this paper.
3.3 Distribution of intracluster ultra compact dwarf galaxies
As mentioned in Sect. 1, Drinkwater et al. (2001) argued based on their higher velocity dispersion that dwarf galaxies might be preferentially infalling. Tracers of disrupted dwarf galaxies are the so-called ultra compact dwarf galaxies (UCDs), the most massive of them being stripped nuclei of formerly nucleated dwarf ellipticals (Hilker et al., 1999; Drinkwater et al., 2003). In Fig. 13 we overlay the density of UCDs from Cantiello et al. (2020) onto the eROSITA image. Overall, we find a broad correlation between the UCD densities and the X-ray surface brightness. In particular, Finger 3 sticks out while the correspondence to the other features is weaker. The elongation in east-west direction in the central part also seems consistent with that observed in Fig. 12.
@@ 3.4 Distribution of intracluster globular clusters
Cantiello et al. (2020) also presented the spatial distribution of globular clusters in the Fornax cluster. The bright globular clusters in the inner region of sq.deg around NGC 1399 have been confirmed spectroscopically as well (Chaturvedi et al., 2022). The globular cluster density is shown in Fig. 14 on top of the eROSITA image. Qualitatively, we find the same as for the UCD distribution – east-west elongation in the center and a correspondence with the X-ray surface brightness in the outer parts. Especially Finger 1 and 2 and Excess 1 and 2 and the Bridge seem to be traced approximately.
@@ 3.5 Distribution of background clusters and superclusters
eROSITA has discovered more than 12,000 galaxy clusters in more than 13,000 sq.deg across the western Galactic hemisphere (Bulbul et al., 2024; Kluge et al., 2024), including the region of the Fornax cluster. Therefore, we can expect about one background cluster per sq.deg, resulting in a large number to lie in the field under study here. As described in Sect. 2, we excised compact sources, which we expect to include background clusters as most of them will be at comparatively high redshift, (the median redshift of the cluster catalog is 0.31), resulting in a small apparent size. Still, if there was an overdensity of clusters in a certain area, either just projected along the line of sight or forming a real supercluster, there might be significant summed emission present from their outskirts or from filaments connecting them. We want to test this in order to avoid incorrectly assigning excess emission to the outskirts of the Fornax cluster. To do so, we overlay in Fig. 15 the eRASS1 background clusters. Their redshifts are labeled and the circle sizes correspond to their radii.
@@
Several things can be noted in this Figure: (i) Our source detection indeed excises almost all of the background clusters, with a region size that is often comparable to . While the source detection process here is very different, this is nonetheless not surprising given that we use data with five times the exposure time compared to eRASS1. It is also not surprising that there is a small fraction of clusters that we do not detect because (a) the purity of the eRASS1 catalog is estimated to be 86% (Bulbul et al., 2024) and (b) a few of them are very nearby () and, therefore, not compact. (ii) Some of the regions of enhanced surface brightness in the field indeed do correspond to overdensities of clusters, sometimes from a few at the same redshift, sometimes at different redshifts. While the clusters themselves are essentially all excised, there appears to be emission well beyond their radii that, combined, is picked up by eROSITA. The cause of this apparent large-scale structure emission (e.g., WHIM surrounding clusters or an overdensity of undetected galaxies, AGN, or small groups of galaxies surrounding the clusters) will be studied in a future paper. (iii) Most of the main excess regions described in previous sections do not correspond to such cluster overdensities, making it unlikely they are due to projection of emission from the outskirts of background clusters. (iv) There are, however, a few X-ray excess regions that do show background cluster overdensities. Particularly interesting is that there are three clusters at the same redshift (0.31) aligned almost perfectly with Finger 1 that points in the north-west direction (projected roughly at the Fornax ). The projected extent is about 15 Mpc, in principle reasonable for a filament. Nonetheless, it seems unlikely that this emission finger is due to a background filament, especially because it would have to be a rather bright and thick (10 Mpc) filament.
@@ 3.6 HI-tail galaxies
Cold gas stripping is widespread for galaxies falling into galaxy clusters (e.g., Cortese et al., 2021). Whether the same processes happen in lower mass galaxy groups is less clear. With a blind ATCA HI survey, Loni et al. (2021) did find evidence for neutral hydrogen removal within the virial region of the Fornax cluster. Using MeerKAT, Serra et al. (2023) recently increased the number of known HI-tail galaxies in Fornax from one to six, showing that ram pressure stripping can operate in this group, especially for galaxies that underwent interactions, thereby making their gas less bound. Four out of these six are located at projected radii beyond , thereby making their X-ray environment inaccessible to XMM-Newton or Chandra.
In Fig. 16, we show the HI contours from Serra et al. (2023) on top of the eROSITA image. None of the host galaxies themselves shows strong X-ray emission. All four galaxies are in a region where X-ray emission from the Fornax ICM is clearly detected by eROSITA, though. Intriguingly, the HI-tails for three of them (FCC 306, NGC 1437B, and NGC 1437A) are parallel to the south-eastern Finger 3 and coincident with an overdensity of possibly infalling UCDs (Fig. 13). The tail of NGC 1427A located inside also points in the same direction. Furthermore, the tails of the two northern galaxies, ESO 359-063 and ESO 358-051 align with the direction towards Excess 4 and also seem coincident with a UCD overdensity.
@@ 3.7 NGC 1316
NGC 1316 is the central galaxy of a galaxy group that appears to be bound to the Fornax cluster and falling into it (Drinkwater et al., 2001). It also hosts Fornax A, a bright radio AGN exhibiting very extended lobes. In Fig. 17, we show the eROSITA X-ray surface brightness and overlaid on it the MeerKAT radio contours from Maccagni et al. (2020). One can notice that the elongation of the X-ray surface brightness aligns with that of the radio lobes. Furthermore, we see in this image that the X-ray emission extends much farther than the radio lobes. The total extent is clearly more than 1 deg. Actually, looking at Fig. 2, bottom, the total extent might well be 2 deg in this direction. This large extent, as well as the fact that the X-ray emission does not trace the radio lobes exactly, shows that the soft X-ray emission traced by eROSITA is dominated by thermal emission from the NGC 1316 intragroup gas instead of by inverse Compton scattering of the photon background on the relativistic electron population that gives rise to the synchrotron emission observed in the radio regime. This appears to be in contradiction to earlier studies that assume that at 1 keV all X-ray emission has a non-thermal origin (e.g., Persic & Rephaeli, 2019). At higher energies 1.5 keV, the non-thermal component has been shown by Suzaku spectral analysis to dominate instead (e.g., Tashiro et al., 2009).
@@ 3.8 SLOW comparison
The SLOW simulations (Sorce, 2018; Dolag et al., 2023) aim to reproduce the local large-scale structure by convolving randomly generated initial conditions with observed large-scale structure data (Sorce et al., 2016). A direct consequence is a set of simulated counterparts for locally observed galaxy clusters (Hernández-Martínez et al., 2024) reproducing positions and general properties with good accuracy. Recent comparisons to observations (e.g., Dietl et al., 2024) have shown especially the large-scale environment components of matched clusters such as massive neighbors and filaments to be matching well due to the reliable constraints on these scales. To test whether this holds for the relatively low-mass Fornax cluster we produce a mock observation of the simulated counterpart of the Fornax cluster in the SLOW simulation. We use the post-processing program SMAC (Dolag et al., 2005) to simulate X-ray emission in the 0.1–2.4 keV band. Additionally, we optimize the observer position for the mock observation so that Fornax is in exactly the correct sky position in order to eliminate projection uncertainties generated by the slight offsets from the true positions the simulation naturally generates.
In Fig. 18, we show the result of the SLOW simulation (Dolag et al., 2023; Hernández-Martínez et al., 2024) at the position of the Fornax cluster. Comparison to Fig. 2 shows, interestingly, that the Bridge towards Fornax A is also present. Furthermore, emission “fingers” to the west can be clearly seen in the image, which resemble those in the eROSITA image. Scrutinizing the simulation data, it turns out that, indeed, the finger to the south-west is co-located with the Fornax cluster. However, the emission in the image corresponding to the finger to the north-west is actually located 53 Mpc behind the Fornax cluster and is due to a projected galaxy group. When looking in detail at the grey gas density contours, though, a weak higher-density extension becomes obvious towards the north-west, right where the eROSITA image shows Excess 1.
@@ 4 Discussion and conclusions
With the new eROSITA observations, we have discovered X-ray emission from the Fornax cluster beyond for the first time. The emission is not symmetric but instead shows preferred directions of higher X-ray surface brightness, as one might expect from gas infall along filaments. E.g., we find finger-like structures with sharp surface brightness drops around , which then slowly fade out to much larger distances as regions of excess emission.
@@
At different levels, the NED member galaxy, UCD galaxy, and globular cluster distributions are all correlated with the X-ray excess regions, and also the HI-tails of some member galaxies point in the same preferred directions. This offers the intriguing possibility that we are witnessing the ongoing growth of the Fornax cluster along preferred infall directions, consistently with the Fornax counterpart in the SLOW simulation. Alternatively, at least some of these features might be related to the ongoing minor merger with NGC 1404. However, we do not find obvious evidence for the shock front predicted by numerical simulations of this merger (Sheardown et al., 2018).
@@
It is worth noting that no perfect agreement among the large-scale structure tracers may necessarily be expected. E.g., for the NED galaxy distribution no distinction is made here between newly infalling galaxies and galaxies that are already in full virial equilibrium with the gravitational potential. The UCD galaxy distribution, on the other hand, might be dominated by infalling galaxies. HI-tails clearly demonstrate infall and the projected direction is clear as well. Last but not least, for the X-ray surface brightness, all corrections have been performed. However, when comparing the surface brightness distribution with the distributions of the other tracers, one needs to be cautious because for none of the tracers based on individual objects (NED galaxies, UCDs, globular clusters, HI-tail galaxies) have any selection effects been taken into account. E.g., as the NED archive is not based on a complete catalog, no selection effect correction of galaxy density estimates can be performed. Therefore, there is uncertainty as different areas might have different sensitivities for the various tracers.
@@
We have furthermore discovered a Bridge of X-ray emission towards NGC 1316, hosting Fornax A. This Bridge approximately coincides with a region of enhanced Faraday depth measured by Anderson et al. (2021), which itself also indicates the presence of a significant number of thermal electrons (and magnetic fields). The X-ray emission of the galaxy group around NGC 1316 itself extends well beyond 1 deg and is aligned with the radio lobes, indicating interaction between the relativistic and thermal electrons.
@@
The detection of enhanced X-ray surface brightness in the outskirts of background clusters in overdense regions of the field implies that with eROSITA we start going beyond the detection of clusters as individual objects and instead see a Universe of continuous large-scale emission – as expected from cosmological simulations.
@@
Based on the hard band (1.0–2.3 keV) image, we are confident that most of the X-ray features are not due to soft X-ray structure in our own Milky Way; still, it remains difficult to exclude with certainty that some of the X-ray excess regions may be due to foreground structure (see, e.g., the extreme cases of emission of the eROSITA Bubbles projected onto the outskirts of the Virgo and Centaurus clusters, McCall et al. 2024; Veronica et al. 2025). Correlations between X-ray excess regions and various other tracers of structure, like member galaxies and intracluster globular clusters, as employed here, do give confidence. Final confirmation of the distance of individual excess regions may need to await high spectral resolution follow-up observations with XRISM (e.g., XRISM Science Team, 2020), HUBS (e.g., Bregman et al., 2023), or even NewAthena (e.g., Nandra et al., 2013; Cruise et al., 2025) in order to pin down the redshifts of the excess regions from their X-ray emission lines.

### v2

@@ 3.1 X-ray surface brightness distribution
The sector 270–360 deg in Fig. 20 shows a bump starting around , likely related to the Fornax A group. In Fig. 9 we then show the surface brightness in a small sector toward and beyond the Fornax A group. It is clear that significant group emission spans at least 1 deg. Moreover, the indication of an apparent emission bridge between the Fornax cluster and the Fornax A group is also confirmed in the surface brightness profile.
@@
To get a better handle on the origin of these surface brightness extensions and excess regions, we compare them to other tracers of the large-scale structure in and around the Fornax cluster in the next sections.
@@ 3.2 Distribution of cluster member galaxies
The most obvious additional tracer of structure is the distribution of galaxies in and around the Fornax cluster. As the Fornax cluster redshift is 0.005, we selected all objects classified as galaxies from the NASA/IPAC Extragalactic Database (NED)444https://ned.ipac.caltech.edu in the image area with a redshift less than 0.01. All of these were assumed to be member galaxies. With this selection we aimed for a high completeness of member galaxies. Since velocity dispersion measurements for the Fornax cluster reach km/s and above (e.g., Drinkwater et al., 2001; Pota et al., 2018), we wanted to ensure to include all galaxies within at least km/s, which corresponds to a range of at least –. Figure 10 shows the member galaxy distribution.
In the central region up to about , the galaxy distribution exhibits an elongation in the east-west direction. On larger scales, well beyond , one can notice a correlation of X-ray surface brightness excess regions and galaxy density. In particular, Fingers 1 and 2 and Excesses 1, 2, and 3 seem to have counterparts in the galaxy distribution. Moreover, the X-ray emission bridge toward the NGC 1316 group also has an excess of galaxies; this is related to the lopsided distribution of Fornax A group dwarf galaxies toward the northeast discussed in Sect. 1 and Su et al. (2021). The underlying reason for this offset between galaxy distribution and Fornax A intragroup gas might be related to the collisional gas lagging behind the collisionless galaxies in their common motion, for example due to ram pressure relative to the ICM of the Fornax cluster outskirts; however, a more detailed investigation of this scenario is beyond the scope of this paper.
3.3 Distribution of intracluster ultra-compact dwarf galaxies
As mentioned in Sect. 1, Drinkwater et al. (2001) argued based on their higher velocity dispersion that dwarf galaxies might be preferentially infalling. Tracers of disrupted dwarf galaxies are the so-called ultra-compact dwarf galaxies (UCDs), the most massive of them being stripped nuclei of formerly nucleated dwarf ellipticals (Hilker et al., 1999; Drinkwater et al., 2003). In Fig. 11 we overlay the density of UCDs from Cantiello et al. (2020) onto the eROSITA image. Overall, we find a broad correlation between the UCD densities and the X-ray surface brightness. In particular, Finger 3 sticks out while the correspondence to the other features is weaker. The elongation in the east-west direction in the central part also seems consistent with that observed in Fig. 10.
@@ 3.4 Distribution of intracluster globular clusters
Cantiello et al. (2020) also presented the spatial distribution of globular clusters in the Fornax cluster. The bright globular clusters in the inner region of sq.deg around NGC 1399 have been confirmed spectroscopically as well (Chaturvedi et al., 2022). The globular cluster density is shown in Fig. 12 on top of the eROSITA image. Qualitatively, we find the same as for the UCD distribution – east-west elongation in the center and a correspondence with the X-ray surface brightness in the outer parts. Especially Fingers 1 (F1) and 2 (F2), Excess 1 (E1), and the bridge seem to be traced approximately.
@@ 3.5 Distribution of background clusters and superclusters
eROSITA has discovered more than 12,000 galaxy clusters in more than 13,000 sq.deg across the western Galactic hemisphere (Bulbul et al., 2024; Kluge et al., 2024), including the region of the Fornax cluster. Therefore, we can expect about one background cluster per square degree, resulting in a large number to lie in the field under study here. As described in Sect. 2, we excised compact sources, which we expected to include background clusters as most of them will be at comparatively high redshifts, (the median redshift of the cluster catalog is 0.31), resulting in a small apparent size. Still, if there was an overdensity of clusters in a certain area, either just projected along the line of sight or forming a real supercluster, there might be significant summed emission present from their un-excised outskirts or from filaments connecting them. We tested this in order to avoid incorrectly assigning excess emission to the outskirts of the Fornax cluster. To do so, we overlay in Fig. 13 the eRASS1 background clusters. Their redshifts are labeled and the circle sizes correspond to their radii.
@@
Several things can be noted in this figure: (i) Our source detection indeed excises almost all of the background clusters, with a region size that is often comparable to . While the source detection process here is very different, this is nonetheless not surprising given that we used data with five times the exposure time compared to eRASS1. It is also not surprising that there is a small fraction of clusters that we do not detect because (a) the purity of the eRASS1 catalog is estimated to be 86% (Bulbul et al., 2024) and (b) a few of them are very nearby () and, therefore, not compact. (ii) Some of the regions of enhanced surface brightness in the field indeed do correspond to overdensities of clusters, sometimes from a few at the same redshift, sometimes at different redshifts. While the clusters themselves are essentially all excised, there appears to be emission well beyond their radii that, combined, is picked up by eROSITA. The cause of this apparent large-scale structure emission (e.g., WHIM surrounding clusters or an overdensity of undetected galaxies, active galactic nuclei, or small groups of galaxies surrounding the clusters) will be studied in a future paper. (iii) Most of the main excess regions described in previous sections do not correspond to such cluster overdensities, making it unlikely they are due to projection of emission from the outskirts of background clusters. (iv) There are, however, a few X-ray-excess regions that do show background cluster overdensities. Particularly interesting is that there are three clusters at the same redshift (0.31) aligned almost perfectly with Finger 1 that points in the northwest direction (projected roughly at the Fornax ). The projected extent is about 15 Mpc, in principle reasonable for a filament (Galárraga-Espinosa et al., 2024). Nonetheless, it seems unlikely that this emission finger is due to a background filament, especially because it would have to be a rather bright and thick (10 Mpc) filament. Indeed, its surface brightness is comparable to that of the few rare filaments individually detected in X-rays (Dietl et al., 2024; Reiprich et al., 2021), which are all at least five times less distant, and the typical width of a filament in numerical simulation is of order 1 Mpc (Galárraga-Espinosa et al., 2021; Wang et al., 2024). Therefore, we conclude that this particular alignment of three projected background clusters with Finger 1 is a coincidence and that their outskirts or filament emission is not responsible for a large fraction of the observed X-ray excess.
In summary, we do find some hints for diffuse emission associated with the outskirts of background clusters but conclude that none of the features discussed here (fingers and excesses) are caused by them.
@@ 3.6 HI-tail galaxies
Cold gas stripping is widespread for galaxies falling into galaxy clusters (e.g., Cortese et al., 2021). Whether the same processes happen in lower-mass galaxy groups is less clear. With a blind ATCA HI survey, Loni et al. (2021) did find evidence of neutral hydrogen removal within the virial region of the Fornax cluster. Using MeerKAT, Serra et al. (2023) recently increased the number of known HI-tail galaxies in Fornax from one to six, showing that ram pressure stripping can operate in this group, especially for galaxies that underwent interactions, thereby making their gas less bound. Four out of these six are located at projected radii beyond , thereby making their X-ray environment inaccessible to XMM-Newton or Chandra.
In Fig. 14 we show the HI contours from Serra et al. (2023) on top of the eROSITA image. None of the host galaxies themselves shows strong X-ray emission. All four galaxies are in a region where X-ray emission from the Fornax ICM is clearly detected by eROSITA, though. Intriguingly, the HI tails for three of them (FCC 306, NGC 1437B, and NGC 1437A) are parallel to the southeastern Finger 3 and coincident with an overdensity of possibly infalling UCDs (Fig. 11). The tail of NGC 1427A located inside also points in the same direction. All of their tails point away from the center of the Fornax cluster, consistent with them falling in for the first time, loosing their HI gas along the way due to ram pressure stripping despite the relatively low ICM density in the Fornax outskirts. Furthermore, the tails of the two northern galaxies, ESO 359-063 and ESO 358-051 align with the direction toward Excess 4 and also seem coincident with a UCD overdensity, although their tails point away from the Fornax center. We conclude that, in the Fornax cluster, HI gas is being stripped from galaxies that are (still) far away from the cluster center and that are moving along preferred infall directions; this might also have implications for ICM heavy element enrichment processes.
@@ 3.7 NGC 1316
NGC 1316 is the central galaxy of a galaxy group that appears to be bound to the Fornax cluster and falling into it (Drinkwater et al., 2001). It also hosts Fornax A, a bright radio active galactic nucleus exhibiting very extended lobes. In Fig. 15 we show the eROSITA X-ray surface brightness and overlaid on it the MeerKAT radio contours from Maccagni et al. (2020). One can notice that the elongation of the X-ray surface brightness aligns with that of the radio lobes. Furthermore, we see in this image that the X-ray emission extends much farther than the radio lobes. The total extent is clearly more than 1 deg. Actually, looking at Fig. 2, bottom, the total extent might well be 2 deg in this direction. This large extent, as well as the fact that the X-ray emission does not trace the radio lobes exactly, shows that the soft X-ray emission traced by eROSITA is dominated by thermal emission from the NGC 1316 intragroup gas instead of by inverse Compton scattering of the photon background on the relativistic electron population that gives rise to the synchrotron emission observed in the radio regime. This appears to be in contradiction to earlier studies that assume that at 1 keV all X-ray emission has a nonthermal origin (e.g., Persic & Rephaeli, 2019). At higher energies 1.5 keV, the nonthermal component has been shown by Suzaku spectral analysis to dominate instead (e.g., Tashiro et al., 2009).
@@ 3.8 SLOW comparison
The simulation of the local Universe (SLOW) simulations (Sorce, 2018; Dolag et al., 2023) aim to reproduce the local large-scale structure by convolving randomly generated initial conditions with observed large-scale structure data (Sorce et al., 2016). A direct consequence is a set of simulated counterparts for locally observed galaxy clusters (Hernández-Martínez et al., 2024) reproducing positions and general properties with good accuracy. Recent comparisons to observations (e.g., Dietl et al., 2024) have especially shown that the large-scale environment components of matched clusters such as massive neighbors and filaments match closely due to the reliable constraints on these scales. To test whether this holds for the relatively low-mass Fornax cluster, we produced a mock observation of the simulated counterpart of the Fornax cluster in the SLOW simulation. We used the post-processing program SMAC (Dolag et al., 2005) to simulate X-ray emission in the 0.1–2.4 keV band. Additionally, we optimized the observer position for the mock observation so that Fornax was in exactly the correct sky position in order to eliminate projection uncertainties generated by the slight offsets from the true positions the simulation naturally generates.
In Fig. 16 we show the result of the SLOW simulation (Dolag et al., 2023; Hernández-Martínez et al., 2024) at the position of the Fornax cluster. Comparison to Fig. 2 shows, interestingly, that the bridge toward Fornax A is also present. Furthermore, emission fingers to the west can be clearly seen in the image, which resemble Fingers 1 and 2 in the eROSITA image. Scrutinizing the simulation data, it turns out that, indeed, the finger to the southwest (Finger 2) is colocated with the Fornax cluster. However, the emission in the image corresponding to Finger 1 is actually located 53 Mpc behind the Fornax cluster and is due to a projected galaxy group. When looking in detail at the black gas density contours, though, a weak higher-density extension becomes obvious toward the northwest, right where the eROSITA image shows Excess 1. The white contours show the stellar density, and the similarity to the observed galaxy and globular cluster density (Figs. 10 and 12, respectively) is striking. First, the same elongation in east-west direction is seen. Second, the offset of the galaxy and gas distributions close to the Fornax A group (discussed in Sect. 3.2) is also present. Analyzing the galaxy velocities in the SLOW simulation shows that those galaxies, as well as those close to Finger 2, are currently moving toward the Fornax cluster center. In conclusion, the agreement between the SLOW simulations and the eROSITA and galaxy observations, on the one hand, lends confidence to the realism of the constrained simulations and, on the other hand, robustly aids the interpretation of the observations.
@@ 4 Discussion and conclusions
With the new eROSITA observations, we have discovered X-ray emission from the Fornax cluster beyond for the first time. In the inner parts, , we detect the large-scale sloshing features already known from previous XMM-Newton mosaic observations. In the outer parts, the emission is not symmetric but instead shows preferred directions of higher X-ray surface brightness, as one might expect from gas infall along filaments. In particular, we find finger-like structures with sharp surface brightness drops around , which then slowly fade out to much larger distances as regions of excess emission.
@@
At different levels, the NED member galaxy, UCD galaxy, and globular cluster distributions are all correlated with the X-ray-excess regions, and the HI tails of some member galaxies also point in the same preferred directions. This raises the intriguing possibility that we are witnessing the ongoing growth of the Fornax cluster along preferred infall directions, a growth that matches that of the Fornax counterpart in the SLOW simulation. The cluster assembly is still ongoing along filaments where preprocessed galaxy groups (with or already without their own X-ray gas) reach the cluster core in specific directions. In the Fornax cluster, this led to an east-west elongation of the galaxy, UCD, and globular cluster distributions (Figs. 10 through 12). In these regions UCDs trace dissolving, nucleated galaxies, with some globular clusters still related in phase space to the accretion event (e.g., Napolitano et al. 2022 and Chaturvedi et al. 2022). Since intracluster globular clusters are also related to the intracluster light and dark matter distribution in galaxy clusters, all these tracers map the mass assembly of the galaxy cluster, as does the X-ray gas (e.g., Forbes et al., 2012; Diego et al., 2023; Martis et al., 2024). All these tracers in the Fornax cluster therefore seem to outline the preferred infall directions. Alternatively, at least some of these features might be related to the ongoing minor merger with NGC 1404. However, we do not find obvious evidence of the shock front predicted by numerical simulations of this merger (Sheardown et al., 2018).
@@
It is worth noting that no perfect agreement among the large-scale structure tracers may necessarily be expected. For example, for the NED galaxy distribution no distinction is made here between newly infalling galaxies and galaxies that are already in full virial equilibrium with the gravitational potential. The UCD galaxy distribution, on the other hand, might be dominated by infalling galaxies. HI tails clearly demonstrate infall, and the projected direction is clear as well. Last but not least, for the X-ray surface brightness, all corrections have been performed. However, when comparing the surface brightness distribution with the distributions of the other tracers, one needs to be cautious because for none of the tracers based on individual objects (NED galaxies, UCDs, globular clusters, or HI-tail galaxies) have any selection effects been taken into account. For example, as the NED archive is not based on a complete catalog, no selection effect correction of galaxy density estimates can be performed. Therefore, there is uncertainty as different areas might have different sensitivities to the various tracers.
@@
We have furthermore discovered a bridge of X-ray emission toward NGC 1316, which hosts Fornax A. This bridge approximately coincides with a region of enhanced Faraday depth as measured by Anderson et al. (2021), which itself also indicates the presence of a significant number of thermal electrons (and magnetic fields). The X-ray emission of the galaxy group around NGC 1316 itself extends well beyond 1 deg and is aligned with the radio lobes, indicating interaction between the relativistic and thermal electrons.
@@
The detection of higher X-ray surface brightness in the outskirts of background clusters in overdense regions of the field implies that with eROSITA we start going beyond the detection of clusters as individual objects and instead see a Universe of continuous large-scale emission – as expected from cosmological simulations.
@@
Based on the hard band image (1.0–2.3 keV), we are confident that most of the X-ray features are not due to soft X-ray structure in our own Milky Way; still, it remains difficult to confidently rule out the possibility that some of the X-ray-excess regions are due to foreground structure (see, e.g., the extreme cases of emission of the eROSITA bubbles projected onto the outskirts of the Virgo and Centaurus clusters; McCall et al. 2024; Veronica et al. 2025). Correlations between X-ray-excess regions and various other tracers of structure, such as member galaxies and intracluster globular clusters, as employed here do support our conclusions. Final confirmation of the distance of individual excess regions may need to await high-spectral-resolution follow-up observations with XRISM (e.g., XRISM Science Team, 2020), HUBS (e.g., Bregman et al., 2023), or even NewAthena (e.g., Nandra et al., 2013; Cruise et al., 2025), from which one will be able to determine the redshifts of the excess regions from their X-ray emission lines.